Write an equation in slope-intercept for: slope -2, passes through (-4,6)
Answer by nyc_function(2741)

Use the point-slope formula.
y - y1 = m(x - x1)
Let m = slope
So, m = -2
Let y1 = 6
Let x1 = -4
y - 6 = -2(x -(-4))
y - 6 = -2(x + 4)
y - 6 = -2x - 8
y = -2x - 8 + 6
y = -2x - 2
Did you follow?